### v1

# Leveraging Distribution Matching to Make Approximate Machine Unlearning Faster

Abstract
Approximate machine unlearning (AMU) enables models to ‘forget’ specific training data through specialized fine-tuning on a retained dataset subset. However, processing this retained subset still dominates computational runtime, while reductions of epochs also remain a challenge. We propose two complementary methods to accelerate classification-oriented AMU. First, Blend, a novel distribution-matching dataset condensation (DC), merges visually similar images with shared blend-weights to significantly reduce the retained set size. It operates with minimal pre-processing overhead and is orders of magnitude faster than state-of-the-art DC methods. Second, our loss-centric method, Accelerated-AMU (A-AMU), augments the unlearning objective to quicken convergence. A-AMU achieves this by combining a steepened primary loss to expedite forgetting with a novel, differentiable regularizer that matches the loss distributions of forgotten and in-distribution unseen data. Our extensive experiments demonstrate that this dual approach of data and loss-centric optimization dramatically reduces end-to-end unlearning latency across both single and multi-round scenarios, all while preserving model utility and privacy. To our knowledge, this is the first work to systematically tackle unlearning efficiency by jointly designing a specialized dataset condensation technique with a dedicated accelerated loss function. Code is available at https://github.com/algebraicdianuj/DC_Unlearning.
Figure 1: The overall pipeline of our proposal to make accelerate AMU, which comprise data-centric and loss-centric components
1 Introduction
Machine unlearning (MU) seeks to expunge specific training points from a deployed model. Exact MU [2, 1] offers perfect deletion guarantees by fully retraining, often via data‐sharding schemes, but remains prohibitively slow for modern networks. Approximate MU (AMU) [13, 15, 6, 9, 25, 11] replaces such guarantees with speed, typically minimizing in iterative or one-step settings:
$\alpha\,\underbrace{\textbf{Main Loss}}_{\text{on retained data}}\,+\,\beta\,\underbrace{\textbf{Forgetting Reg.}}_{\text{drives deletion}}$
Here, $\alpha,\beta\geq 0$ control the utility-vs-forgetting trade-off. We focus on the common iterative case $\alpha=1$, as one step variants assume unrealistically simple loss landscapes. Although approximate machine unlearning (AMU) avoids full retraining, repeatedly scanning the large retained set still dominates runtime. At the same time, reducing unlearning epochs remains a challenge, analogous to how BatchNorm accelerated training convergence [12].
We address this by realizing reduction of the retained dataset via specialized speed-focused dataset condensation [30]. Building upon distribution-matching dataset condensation [29], we propose Blend [^1] This method is a revision of our previous work available at https://arxiv.org/abs/2402.00195. , that learns a compact blend-weight vector to synthesize a tiny set of synthetic datasets that preserve training utility. Integrating this into AMU pipeline shrinks the retained set and accelerates unlearning without sacrificing accuracy or privacy, with a minor overhead. We further push for loss-centric optimization through Accelerated-AMU (A-AMU) loss that steepens the main loss landscape, alongside loss distribution matching formulation (as a MIA minimizer), enabling rapid convergence to the unlearned model. Our approach demonstrates substantial speedups and very strong unlearning compared to baseline SOTA unlearning methods. We summarize our contributions as follows:
  • Blend DC, specifically designed for unlearning, combines image blending with distribution matching, achieving an average 44% retained dataset reduction with approximately 7.5% preprocessing overhead compared to baseline unlearning, and is approximately 1500 times faster than state-of-the-art dataset condensation methods.
  • A-AMU augments the AMU loss formulation with a steepened primary loss and a differentiable membership-inference regularizer, reducing unlearning runtime by 76.82% in single-round and 51.83% in multi-round scenarios compared to baseline methods, by minimizing the number of epochs, while maintaining strong privacy-utility statistics.
  • We benchmarked our approaches on various image classification datasets and models, achieving 84.61% and 54.45% faster end-to-end unlearning in single and multi-round scenarios, respectively, effectively filling latency gaps compared to current state-of-the-art AMU methods.
2 Related Works
Machine unlearning (MU) has evolved along two complementary axes. Exact MU guarantees equivalence to full retraining by either retraining on data shards (SISA [1], ARCANE [27]) or deriving closed-form updates for simple models such as $k$-means and linear classifiers [8, 19], but scaling these ideas to over-parameterized deep networks remains hard. Approximate MU relaxes this guarantee for speed: influence-function or Hessian sketches explicitly subtract a point’s contribution [10, 26, 9, 25];“bad-teacher” distillation removes the target data then distills utility back from a clean model [15, 6]; and recent accelerations prune parameters [21, 13], nudge decision boundaries [5], or couple sparse adapters with shard partitioning, culminating in class-agnostic, data-free DELETE [32] and Unlearn-and-Distill [16]. Parallel to MU, dataset condensation (DC) compresses training sets into compact proxies: from gradient matching [24, 30] and distribution/trajectory matching [29, 4] to latent re-parameterization, differentiable augmentation and infinite-width kernels [14, 28, 20], with EDF recently highlighting class-discriminative cues via Grad-CAM [23]. However DC for MU is nascent: QuickDrop naively integrates gradient-matching DC into federated unlearning [7], and TCGU shows graph-level gains by naively condensing once then editing the proxy [17]; nevertheless, existing DC objectives are still too costly to serve as a drop-in unlearning accelerator, motivating our targetted and simpler, distribution-matching approach.
3 Methodology
3.1 Preliminaries
Let the training dataset $\mathbf{D}$, consist of labeled images, denoted as $\{I_{p}\}_{p}$, comprising of $C$ classes and $N_{c}$ images per class, where $c\in\{1,2,\dots,C\}$. The training dataset be decomposed disjointly into retain $\mathbf{R}$ and forget $\mathbf{F}$ dataset. $\mathbf{D}=\{I_{p}\}_{p}$ was used to train a parameterized target network, and it achieved local minima parameters $\theta$ on average loss $\mathcal{L}(\theta;x,y)$ over $(x,y)\in\mathbf{R}\cup\mathbf{F}$. Let $\mathbf{D}$ and $\mathbf{\tilde{T}}$ be datasets drawn from the same underlying distribution $\mathcal{D}$, where $\mathbf{\tilde{T}}$ is the testing dataset, and is used in evaluation of training. Our objective is to unlearn forget set $\mathbf{F}\subset\{I_{p}\}_{p}$ from the pretrained model. Denote by $\mathcal{D}_{R}$ and $\mathcal{D}_{F}$ the empirical distribution of $\mathbf{R}$ and $\mathbf{F}$ respectively. The subset $\mathbf{T}\subseteq\mathbf{\tilde{T}}$ consists of elements whose class labels match those in $\mathbf{F}$, and $\mathcal{D}_{T}$ is its empirical distribution.
3.2 Blend Dataset Condensation
Under Algorithm 1, we divide images of each class of D into $k$ sub-classes using $k$-means clustering over the feature representations, extracted by a randomly initialized lightweight feature extractor $\psi_{\vartheta}$ (much smaller than the pretrained model to be unlearned). The resulting partition of the dataset can be written as:
[…]
where $i$ indexes the sub-class (with a total of $k\times C$ sub-classes), and $j$ indexes the images within each sub-class. Each $i$-th subclass contain $n_{i}$ images. For each sub-class $i$, we define a blended image $\mathcal{I}(\omega_{i})$ as a weighted average of the images $\{I_{ij}\}_{j=1}^{n_{i}}$ b elonging to that sub-class:
$\mathcal{I}(\omega_{i})\overset{\mathrm{blend}}{=}\frac{1}{\sum_{j=1}^{n_{i}}\omega_{ij}}\sum_{j=1}^{n_{i}}\omega_{ij}\,I_{ij}$ (2)
Algorithm 1 $k$-means Partitioning ($\mathbf{D}=\{I_{p}\}_{p}$)
1:Indices: $c=1,\dots,C$ (class); $\kappa=1,\dots,k$ ($k$-means cluster); $b=1,\dots,n_{\kappa}$ (images in cluster $\kappa$)
2:$\mathcal{C}\leftarrow\varnothing$triangleright Initialize image clusters
3:for each class $c=1,\dots,C$ do
4:    $\{I_{a}\}_{a=1}^{N_{c}}\leftarrow\{I_{p}\mid y_{p}=c\}$triangleright Collect class images
5:    $\{\psi_{\vartheta}(I_{a})\}_{a=1}^{N_{c}}$triangleright Compute feature vectors
6:    $\{\psi_{\vartheta}(I_{a})\}_{a=1}^{N_{c}}\xrightarrow{k\text{-means}}\bigl{\{}\{\psi_{\vartheta}(I_{b\kappa})\}_{b=1}^{n_{\kappa}}\bigr{\}}_{\kappa=1}^{k}$triangleright Run $k$-means
7:    $\bigl{\{}\{\psi_{\vartheta}(I_{b\kappa})\}_{b=1}^{n_{\kappa}}\bigr{\}}_{\kappa=1}^{k}\to\bigl{\{}\{I_{b\kappa}\}_{b=1}^{n_{\kappa}}\bigr{\}}_{\kappa=1}^{k}$triangleright Map to images
8:    $\mathcal{C}\leftarrow\mathcal{C}\cup\bigl{\{}\{I_{b\kappa}\}_{b=1}^{n_{\kappa}}\bigr{\}}_{\kappa=1}^{k}$triangleright Add to global set
9:end for
10:$\mathcal{C}=\bigl{\{}\{\{I_{b\kappa c}\}_{b=1}^{n_{\kappa c}}\}_{\kappa=1}^{k}\bigr{\}}_{c=1}^{C}$triangleright Full cluster set
11:Refine notation: $b\mapsto i$, $(\kappa,c)\mapsto j$
12:return $\{\{I_{ij}\}_{i=1}^{n_{j}}\}_{j=1}^{k\times C}$triangleright Partitioned D dataset
where $\omega_{ij}$ are scalar weights associated with the $j$-th image in sub-class $i$. These weights are treated as learnable parameters and are optimized to ensure that the feature representation of the blended image $\mathcal{I}_{i}$ closely matches the mean feature representation of its constituent images to form a condensed single image proxy of $\{I_{ij}\}_{j=1}^{n_{i}}$. We estimate $\omega_{i}$ in equation (2), via following distribution matching [29] loss:
$\min_{\omega_{i}}\;\mathbb{E}_{\vartheta\sim P_{\vartheta}}\Bigl{\|}\tfrac{1}{n_{i}}\sum_{j=1}^{n_{i}}\psi_{\vartheta}(I_{ij})\;-\;\psi_{\vartheta}\bigl{(}\mathcal{I}(\omega_{i})\bigr{)}\Bigr{\|}_{2}^{2}$
[…]
3.3 Creating the Reduced Retain Dataset
Dataset condensation replaces each original feature $x\!\in\!\mathbb{R}^{n}$ (as Lipchitz mapping of original image) with a synthetic prototype whose placement preserves the topology of the manifold on which the features live. Because this process is global, a prototype can still migrate into the neighbourhood of the to-be forgotten feature samples, thus creating an obstacle for unlearning. Furthermore, since the condensation time scales with the size of the retain set, condensing the entire retain dataset becomes a significant computational bottleneck. We therefore propose a targeted dataset condensation on specific partition of retain dataset, whose features are most far away from those of forget samples. To achieve this, we propose following steps:
Step 1: Partition the retain set with sampling $\mathcal{F}$
Algorithm 2 systematically separates the retain images (using partitioned training set and $\mathbf{F}$) into:
  • free images $\mathbf{R}_{\text{free}}$ whose features are distant from $\mathbf{F}$
  • residual images $\mathbf{R}_{\text{residual}}$ that lie close to features of  $\mathbf{F}$
Specifically, the inter-cluster indicator $\phi(j)$ and the intra-cluster indicator $\Phi(i)$ test whether the set of images $\{I_{ij}\}_{j=1}^{n_{i}}$ for each cluster $j$ intersects the forget set $\mathbf{F}$, thereby partitioning $R$ into $\mathbf{R}_{\text{free}}$ and $\mathbf{R}_{\text{residual}}$.
Step 2: Condense only the free retain subset
Each cluster in the free retain subset (i.e. those with $\phi(j)\neq-1$) is condensed by blending its $n_{j}$ images into a single prototype, thus creating condensed version of free retain set:
$\mathbf{Condense}\bigl{(}\mathbf{R}_{\text{free}}\bigr{)}\;=\;\Bigl{\{}\{\,I_{ij}\,\}_{j=1}^{n_{i}}\;\xrightarrow{\;\mathrm{blend}\;}\mathcal{I}\!\bigl{(}\omega_{i}\bigr{)}\Bigr{\}}_{i=\phi(1)}^{\phi(kC)}.$ (3)
allowing no synthetic point converging towards  $\mathbf{F}$ features.
Step 3: Create the Reduced Retain Dataset
Having partitioned the original retain set into the distant “free” clusters and the forget images sensitive “residual” images, and condensed only the former via the blending operation in Eq. (3), we now assemble our final training set.
[…]
This reduced retain dataset preserves all samples that lie near the forget manifold (ensuring precise boundary control during unlearning), while summarising only those clusters that have no intersection with $\mathbf{F}$. As a result, $\mathbf{R}_{\text{reduced}}$ offers the computational savings from total dataset condensation, with advantage of not sacrificing the fine‐grained feature information needed to effectively forget and persist generalization, which would have been possible with raw AMU.
Table 1: Unlearning ResNet-18/CIFAR-10 (mean $\pm$ 95% CI) with $k$-means partitioning time $5.5753\pm 0.1391$ s, $\mathcal{F}$-sampling time $0.1184\pm 0.0107$ s
MIA Score $\downarrow$ Retain Accuracy $\uparrow$ Forget Accuracy $\downarrow$ Test Accuracy $\uparrow$ Unlearning Time (s) $\downarrow$
Method w/o DC w DC w/o DC w DC w/o DC w DC w/o DC w DC w/o DC w DC
Random Class Forgetting (Condensation Time for 1000 IPC: $47.7812\pm 1.3323$ s)
Retraining $49.76\pm 2.57$ $50.51\pm 3.55$ $100.00\pm 0.00$ $75.50\pm 5.63$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $80.35\pm 2.08$ $62.47\pm 4.88$ $1141.71\pm 51.30$ $224.50\pm 0.14$
CF $50.18\pm 1.16$ $55.39\pm 2.67$ $100.00\pm 0.00$ $96.39\pm 0.75$ $00.00\pm 0$ 0.00 $0.30\pm 0.45$ $81.19\pm 2.40$ $77.75\pm 1.25$ $577.81\pm 55.36$ $112.32\pm 0.04$
Bad Distillation $62.83\pm 9.14$ $57.31\pm 7.01$ $96.26\pm 0.72$ $98.88\pm 0.18$ $1.23\pm 3.90$ $24.99\pm 79.51$ $77.33\pm 2.37$ $81.19\pm 6.28$ $284.54\pm 28.72$ $117.04\pm 0.05$
$L_{1}$ Sparse $47.32\pm 2.73$ $53.79\pm 0.88$ $99.00\pm 0.97$ $96.36\pm 0.55$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $78.42\pm 3.40$ $77.66\pm 1.32$ $358.62\pm 28.80$ $69.88\pm 0.01$
SCRUB $50.12\pm 0.37$ $50.00\pm 0.00$ $15.28\pm 7.66$ $11.11\pm 0.00$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $13.80\pm 6.90$ $10.00\pm 0.00$ $241.23\pm 23.22$ $61.58\pm 0.02$
Pruning $50.87\pm 3.22$ $50.80\pm 2.27$ $98.08\pm 0.73$ $100.00\pm 0.00$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $77.43\pm 2.41$ $72.65\pm 1.27$ $235.50\pm 6.20$ $46.74\pm 0.01$
A-CF $51.71\pm 2.29$ $48.25\pm 7.31$ $97.88\pm 4.39$ $98.92\pm 0.62$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $77.01\pm 3.20$ $77.45\pm 3.63$ $75.76\pm 0.27$ $16.01\pm 0.96$
10% Uniform Forgetting (Condensation Time for 1000 IPC: $34.1780\pm 0.7403$ s)
Retraining $50.47\pm 0.61$ $50.05\pm 1.64$ $100.00\pm 0.00$ $91.48\pm 0.41$ $87.48\pm 1.01$ $82.25\pm 1.01$ $86.84\pm 0.62$ $80.65\pm 0.48$ $1132.22\pm 33.71$ $664.73\pm 6.16$
CF $52.45\pm 1.18$ $54.39\pm 0.89$ $100.00\pm 0.00$ $96.95\pm 0.17$ $93.30\pm 0.53$ $93.42\pm 0.08$ $88.11\pm 0.25$ $86.60\pm 0.20$ $561.76\pm 4.92$ $332.23\pm 2.78$
Bad Distillation $82.92\pm 1.51$ $85.78\pm 2.43$ $93.78\pm 0.64$ $94.43\pm 0.98$ $27.99\pm 9.74$ $18.71\pm 9.63$ $78.77\pm 0.56$ $79.57\pm 1.42$ $275.64\pm 0.56$ $264.16\pm 1.54$
$L_{1}$ Sparse $50.87\pm 1.39$ $52.22\pm 0.99$ $98.87\pm 0.23$ $92.81\pm 0.45$ $86.18\pm 0.60$ $86.30\pm 1.07$ $84.36\pm 0.35$ $82.10\pm 0.56$ $348.20\pm 0.98$ $206.46\pm 1.54$
SCRUB $51.27\pm 0.42$ $51.59\pm 1.46$ $99.96\pm 0.01$ $77.94\pm 42.33$ $88.35\pm 1.39$ $73.11\pm 37.70$ $85.99\pm 1.09$ $70.56\pm 34.47$ $235.68\pm 1.63$ $149.79\pm 0.95$
Pruning $51.08\pm 0.72$ $51.75\pm 0.68$ $96.76\pm 0.87$ $97.52\pm 1.28$ $85.67\pm 0.99$ $83.22\pm 2.08$ $82.51\pm 0.70$ $79.75\pm 2.08$ $234.11\pm 0.35$ $138.16\pm 0.96$
A-CF $51.86\pm 0.56$ $51.95\pm 0.47$ $99.53\pm 0.85$ $99.76\pm 0.10$ $86.19\pm 1.41$ $85.42\pm 0.15$ $86.18\pm 2.50$ $86.23\pm 0.38$ $76.24\pm 1.13$ $47.32\pm 3.05$
Algorithm 2 Sampling $\mathcal{F}\!\bigl{(}\mathbf{F},\{\{I_{ij}\}_{i=1}^{n_{j}}\}_{j=1}^{k\times C}\bigr{)}$
1:Indices: $j=1,\dots,kC$ (cluster); $i=1,\dots,n_{j}$ (image in cluster $j$)
2:$\phi(j)=\begin{cases}-1,&\mathbf{F}\cap\{I_{ij}\}_{j=1}^{k\times C}=\emptyset\\ j,&\text{otherwise}\end{cases}$triangleright indexing not $\mathbf{F}$ clusters
3:$\Phi(i)=\begin{cases}i,&\mathbf{F}\cap\{I_{ij}\}_{i=1}^{n_{j}}\\ -1,&\text{otherwise}\end{cases}$triangleright indexing within $\mathbf{F}$ clusters
4:$\mathbf{R}_{\text{free}}=\{\{I_{ij}\}_{i=1}^{n_{j}}\}_{j=\phi(1)}^{\phi(kC)}$triangleright $\mathbf{R}$ subset dissimilar to $\mathbf{F}$
5:$\mathbf{R}_{\text{residual}}=\{\{I_{ij}\}_{i=\Phi_{j}(1)}^{\Phi_{j}(n_{j})}\}_{j=1}^{kC}$triangleright $\mathbf{R}$ subset similar to $\mathbf{F}$
6:return $\mathbf{R}_{\text{free}},\mathbf{R}_{\text{residual}}$
3.4 Accelerated Approximate Unlearning
Carlini et al. [3] report that per-sample classification losses of deep networks are strongly non-Gaussian. They considered to “Gaussianise” losses by ad-hoc monotone maps (e.g. combination of $\exp$ and $\log$ maps), without guaranteeing that the output is normally distributed. We instead apply the probability–integral transform (PIT): for any continuous variable $X$ with CDF $F_{X}$, $Z=\Phi^{-1}\!\bigl{(}F_{X}(X)\bigr{)}$ is standard normal. Substituting $F_{X}$ by a differentiable empirical surrogate yields an end-to-end differentiable pipeline whose transformed losses converge to $\mathcal{N}(0,1)$. Following this motivation, we intend to transform arbitrary random samples, which may not be following Gaussian distribution towards samples following standard normal distribution. Let $x=(x_{1},\ldots,x_{N})\;\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\;F$, such that $F$ is an unknown continuous CDF. The empirical CDF at $x_{i}$ is $\hat{F}_{N}(x_{i})=\tfrac{1}{N}\sum_{j=1}^{N}\mathbf{1}\{x_{j}\leq x_{i}\}.$ Replacing the indicator by the temperature–controlled sigmoid $\sigma_{K}(z)=(1+e^{-Kz})^{-1},\;K>0,$ gives the differentiable estimator
$Q_{i}^{(K)}(x)$ $=\frac{1}{N}\sum_{j=1}^{N}\sigma_{K}\!\bigl{(}x_{i}-x_{j}\bigr{)},\qquad i=1,\ldots,N,$
such that $\lim_{K\to\infty}Q_{i}^{(K)}(x)=\hat{F}_{N}(x)$. We stack all samples into coordinates of vector as
$\mathbf{Q}^{(K)}(x)=\bigl{(}Q_{1}^{(K)}(x),\ldots,Q_{N}^{(K)}(x)\bigr{)}^{\!\top}\in(0,1)^{N}.$
Because each entry of $\mathbf{Q}^{(K)}(x)$ lies in $(0,1)$, apply the probit (quantile) function $\Phi^{-1}$ of the standard normal distribution $\mathcal{N}(0,1)$ (mean 00, variance $1$) to obtain [22]
$\mathbf{z}^{(K)}=\Phi^{-1}\!\bigl{(}\mathbf{Q}^{(K)}(x)\bigr{)}\in\mathbb{R}^{N}.$ (4)
Thus any sample $x$ from a continuous distribution is converted to a sample whose coordinates converge to i.i.d. standard normals as $K\to\infty$: $\mathbf{z}^{(K)}\;\xrightarrow{K\to\infty}\;\Phi^{-1}\!\bigl{(}\hat{F}_{N}(x)\bigr{)}.$ In practice one chooses $K$ finite but large to balance smooth gradients (for optimization) with closeness to the empirical CDF.
3.4.1 Transforming losses of model to standard normal
For the frozen pretrained network with parameters $\theta$, we define the per-sample log transformed network’s original loss (e.g. cross-entropy) as $\ell(x,y)=\log\!\bigl{(}1+\mathcal{L}(\theta;x,y)\bigr{)}$.
Collect all losses of $\mathbf{F}$ and $\mathbf{T}$:
$\bm{\ell}_{F}$ $=\bigl{(}\ell(x_{1},y_{1}),\dots,\ell(x_{n_{F}},y_{n_{F}})\bigr{)}^{\!\top},$
$\bm{\ell}_{T}$ $=\bigl{(}\ell(x^{\prime}_{1},y^{\prime}_{1}),\dots,\ell(x^{\prime}_{n_{T}},y^{\prime}_{n_{T}})\bigr{)}^{\!\top}.$
with $(x_{i},y_{i})\!\sim\!\mathcal{D}_{F}$ and $(x^{\prime}_{i},y^{\prime}_{i})\!\sim\!\mathcal{D}_{T}$.
Applying the vector transformation in (4) to the loss vectors yields $\mathbf{Z}_{F}^{(K)}=\Phi^{-1}\!\bigl{(}\mathbf{Q}^{(K)}(\bm{\ell}_{F})\bigr{)},\qquad\mathbf{Z}_{T}^{(K)}=\Phi^{-1}\!\bigl{(}\mathbf{Q}^{(K)}(\bm{\ell}_{T})\bigr{)},$ where $\mathbf{Q}^{(K)}$ is the differentiable CDF estimator from Section 4.1 and $\Phi^{-1}$ is the probit of the standard normal distribution $\mathcal{N}(0,1)$. As $K\to\infty$ the mapping converges to the empirical probit, and this allows use to play with transformed losses in distribution matching manner. As emphasize previously, practice we keep $K$ sufficiently large to preserve differentiability.
Table 2: Unlearning on ResNet-50/CIFAR-10 (mean $\pm$ 95% CI) with $k$-means partitioning time $5.5753\pm 0.1391$ s, $\mathcal{F}$-sampling time $0.1184\pm 0.0107$ s
MIA Score $\downarrow$ Retain Accuracy $\uparrow$ Forget Accuracy $\downarrow$ Test Accuracy $\uparrow$ Unlearning Time (s) $\downarrow$
Method w/o DC w DC w/o DC w DC w/o DC w DC w/o DC w DC w/o DC w DC
Random Class Forgetting (Condensation Time for 1000 IPC: $47.7812\pm 1.3323$ s)
Retraining $48.61\pm 1.85$ $49.49\pm 2.38$ $100.00\pm 0.00$ $68.45\pm 2.51$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $79.96\pm 1.10$ $56.46\pm 2.81$ $3580.71\pm 94.46$ $705.82\pm 0.30$
CF $51.07\pm 3.71$ $57.87\pm 2.74$ $100.00\pm 0.00$ $97.89\pm 0.18$ $00.00\pm 0$ 0.00 $1.65\pm 1.73$ $81.28\pm 1.33$ $79.11\pm 1.89$ $1789.40\pm 47.15$ $352.38\pm 0.74$
Bad Distillation $57.97\pm 5.13$ $62.99\pm 6.81$ $97.31\pm 1.59$ $99.69\pm 0.09$ $00.00\pm 0$ 0.00 $0.01\pm 0.03$ $77.68\pm 2.36$ $80.64\pm 1.16$ $1370.90\pm 34.32$ $572.67\pm 0.69$
$L_{1}$ Sparse $49.18\pm 2.20$ $58.95\pm 4.72$ $98.52\pm 0.55$ $98.54\pm 0.32$ $00.00\pm 0$ 0.00 $4.03\pm 3.58$ $77.61\pm 2.12$ $79.61\pm 1.49$ $1446.43\pm 6.34$ $288.65\pm 0.31$
SCRUB $50.00\pm 0.00$ $50.00\pm 0.00$ $15.50\pm 3.00$ $11.14\pm 0.06$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $13.88\pm 2.72$ $10.05\pm 0.09$ $759.90\pm 9.19$ $198.01\pm 0.21$
Pruning $48.37\pm 2.43$ $50.56\pm 0.81$ $97.10\pm 1.14$ $99.08\pm 0.63$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $76.85\pm 1.84$ $69.80\pm 3.61$ $722.06\pm 5.91$ $143.73\pm 0.23$
A-CF $52.10\pm 1.50$ $50.99\pm 0.67$ $97.75\pm 1.50$ $95.90\pm 2.73$ $00.00\pm 0$ 0.00 $0.10\pm 0.19$ $77.03\pm 1.08$ $75.90\pm 3.82$ $218.30\pm 0.26$ $44.54\pm 1.80$
10% Uniform Forgetting (Condensation Time for 1000 IPC: $34.1780\pm 0.7403$ s)
Retraining $49.46\pm 0.87$ $49.46\pm 1.29$ $100.00\pm 0.00$ $91.26\pm 0.59$ $87.41\pm 1.05$ $80.88\pm 0.78$ $87.01\pm 0.16$ $80.41\pm 0.64$ $3554.87\pm 92.12$ $2095.34\pm 29.30$
CF $53.32\pm 0.90$ $54.20\pm 0.56$ $100.00\pm 0.00$ $97.79\pm 0.22$ $94.34\pm 0.89$ $95.19\pm 0.36$ $88.64\pm 0.15$ $87.52\pm 0.83$ $1770.13\pm 13.17$ $1046.32\pm 16.59$
Bad Distillation $84.07\pm 1.69$ $88.42\pm 1.62$ $94.77\pm 0.30$ $96.16\pm 0.92$ $23.76\pm 9.60$ $11.93\pm 1.87$ $79.80\pm 1.24$ $81.67\pm 1.86$ $1371.63\pm 35.81$ $1297.61\pm 15.55$
$L_{1}$ Sparse $52.10\pm 1.60$ $53.25\pm 0.38$ $98.25\pm 0.39$ $95.31\pm 0.91$ $87.41\pm 1.22$ $90.13\pm 1.93$ $84.25\pm 1.17$ $84.07\pm 0.90$ $1463.85\pm 38.41$ $856.47\pm 13.84$
SCRUB $51.97\pm 0.91$ $52.32\pm 2.80$ $97.99\pm 3.40$ $86.19\pm 33.28$ $87.38\pm 6.03$ $83.37\pm 30.90$ $84.77\pm 3.59$ $78.17\pm 27.14$ $756.68\pm 1.36$ $483.13\pm 5.28$
Pruning $51.76\pm 0.54$ $51.43\pm 0.93$ $96.81\pm 1.42$ $97.62\pm 0.39$ $86.66\pm 0.69$ $84.71\pm 0.77$ $83.40\pm 1.58$ $81.61\pm 0.68$ $730.16\pm 20.13$ $426.54\pm 6.89$
A-CF $50.20\pm 2.65$ $52.27\pm 1.46$ $95.49\pm 4.84$ $97.06\pm 2.01$ $82.57\pm 3.42$ $84.83\pm 2.89$ $81.72\pm 5.24$